# On the Partitioning of GPU Power among Multi-Instances

### v1

@@ III-B Analysis of Tensor Workloads
Having demonstrated the heterogeneity even within CUDA kernels accomplishing the same end task, we now analyze TENSOR kernels. Fig. 8 shows the distribution of key metrics of POWER, TENSO, and DRAMA across the two workloads of LLM Inference with Llama-3B model (LLAMA) and GPUBurn performing matrix multiplications, both on A100. The power consumption during GPUBurn is higher, with a broader distribution, reflecting the benchmark workload’s impact on power. Similarly, DRAMA is also higher for GPUBurn on average as opposed to LLAMA, however comprising mild peaks at $\sim$0.4. The TENSO activity on LLAMA is also lower on average as compared to GPUBurn. These differences between the two sets of distributions highlight the lack of generalization across workloads with each workload having its own set of characteristics.
Fig. 9 demonstrates the variation in power with TENSO and DRAMA across heterogeneous workloads depicting the wide variation with respect to the range of utilization of the various units as well as the resulting power consumption, compounding the difficulty in unifying power prediction models with lightweight metrics. The challenge is further exacerbated at runtime by the variable durations of workloads. Each workload may run for a different amount of time, resulting in a disproportionate amount of data collected for certain periods while other periods may be underrepresented. This imbalance renders it more difficult to accurately model the system’s behavior across all possible workload conditions. Consequently, the variability renders having a universal model towards power prediction a significant predicament, particularly with the black-box visibility into workloads in execution.
@@ III-C Hardware Heterogeneity
Dynamic variability of runtime monitored metrics on heterogeneous GPU hardware stems from disparities in monitoring granularity and resource allocation visibility across instances. While some workloads exhibit distinct metric patterns due to high resource utilization, others are underrepresented, leading to uneven data distributions. This imbalance poses challenges for constructing robust predictive models, as the variability in metric fidelity fails to capture the nuanced interactions between workloads and GPU resources. Consequently, achieving consistent power prediction accuracy across diverse workload profiles remains a complex task, necessitating tailored approaches to handle such heterogeneity at runtime with evolving workload behavior.
@@ III-D Implications of Metrics
While DCGM metrics are lightweight and easily available, NVIDIA Nsight Compute (NCU) metrics provide more advanced statistical measures into the working of the GPU components. However, obtaining such metrics involve large overheads that makes it impossible to aquire them in production systems. In order to quantify the benefits provided by NCU metrics on the power consumption specifically, we train four models, Linear Regression (LR), Gradient Boosting (GB), Random Forest and XGBoost on the MATMUL workload. We chose the NCU metrics with non-trivial correlation with power for inclusion in the models. DRAM and cache access related metrics showed the highest correlation. Fig. 13 demonstrates the benefit of incorporating NCU metrics versus only relying on DCGM metrics. When using only NCU metrics, there is no significant improvement, indicating that these metrics alone may not provide substantial additional predictive power. The combined graph integrating both DCGM and NCU metrics, however shows significant improvement, suggesting that the inclusion of NCU metrics offers benefits beyond what is achieved with DCGM metrics alone, producing a much lower median error. However, NCU metrics add significant overhead and are difficult to get in production systems. Furthermore, more studies are required to ascertain their value in more complex workloads using tensor cores. So we leverage lightweight DCGM metrics towards power attribution in partitioned GPUs.
Figure 13: DCGM vs NSIGHT compute metrics
@@ III-E Full GPU Power Models
Fig. 14 presents CDF (Cumulative Distribution Function) plots showing error percentages for models trained and tested across different datasets. We evaluate and compare LR, GB, RF, and XGB models. For the Granite Train/Llama Test plot, all four models demonstrate similar performance, reflecting their ability to generalize across workloads with shared characteristics. When training and testing on the same dataset (e.g., Granite Train/Granite Test or Llama Train/Llama Test), errors are low, highlighting the strong fit of the models to consistent data sources. These scenarios reveal minimal variability, demonstrating how well the models adapt to familiar patterns. In contrast, cross-validation across heterogeneous workloads, such as Granite Train/MATMUL Test, results in higher prediction errors. The largest errors occur in the Llama Train/MATMUL Test scenario, reflecting the challenges models face when predicting power consumption for workloads vastly different from the training data. LR struggles the most in these cases, with error percentages often exceeding $\sim$90%, while GB and RF fare slightly better but still exhibit variability. Cross-validation tests further underline the importance of training data diversity. Models trained on granite and tested on GPUBurn or MATMUL show a clear increase in error percentages, particularly for LR and GB.
To improve generalization, it is crucial to curate representative training sets that encompass a wide range of workload behaviors. This approach helps mitigate errors and enhances the models’ ability to handle diverse scenarios effectively, as demonstrated by the unified model depicted in the correlation matrix as in Fig. 15. However, the variability across workloads and predictor models compound the difficulty of a unified offline model spanning across the vast diversified workloads. Thus, unified workloads trained on observed runtime behaviour can abate this conundrum. A unified workload on runtime workloads capture local hardware behaviour coupled with workload specific characteristics towards accurate power prediction. A unified model trained on combined datasets (granite, llama, MATMUL, and GPUburn) achieves the lowest error percentages on average. This indicates that training on aggregated data enables the model to capture broader workload behaviors, unlike models trained on isolated datasets. LR, however, consistently underperforms due to its inability to capture non-linear GPU metric-power relationships, while RF emerges as the most robust predictor on average. This depicts that there is no one-methodology-fits-all for predicting the power consumption with respect to heterogeneous workloads.
Figure 15: Correlation Matrix for Individual Workloads and the Unified data
@@ IV Partitioning GPU Power
In this section, using GPU power characteristics learned through benchmarking, we design approaches for partitioning the aggregate GPU power among its MIG partitions. This approach can be adapted with minimal changes to other forms of shared execution such as time-sharing and multi-process service (MPS) also. The high-level principles should be applicable in general to GPUs from other manufacturers such as Intel and AMD.
@@
Compute processors, including GPUs, are known to suffer from lack of energy proportionality[27], and consume power even in idle state when no process is active. Such power is referred to as idle power, which is generally a non-trivial amount. The idle power of NVIDIA A100-80GB is $\sim$85 W at frequencies above 1200 MHz against its peak power of $\sim$400-450 W. When a GPU executes some process, additional power is consumed, which is referred to as the active power. A GPU’s total power is the sum of its idle and active powers.
@@ IV-A Estimation via Generic Full GPU Power Models
This method uses the unified model (which is constructed using training data from multiple representative workloads) as described earlier to estimate MIG instance power. To begin with, the unified model is used to predict the power consumption of each MIG instance individually using the instance’s normalized DCGM metrics as features. Thus, each MIG prediction is for the power that would be consumed when just that MIG’s workload executes on the GPU, and hence would include the full GPU idle power. The full GPU idle power is deducted from the prediction to get the MIG’s active power. Next, idle power for instance $k$ is computed as described at the beginning of the section. For illustration, suppose that only 2G and 3G partitions are active, then they will be assigned $\frac{2}{5}^{\rm th}$ and $\frac{3}{5}^{\rm th}$ of the idle power, resp. Finally, the total power for an instance is obtained by adding its idle and active powers. The accuracy of the overall approach is evaluated by summing all instance powers and comparing the sum to the total measured GPU power.
@@
TABLE II: MIG Instance combinations for evaluating power partitioning.
Tab. II shows the various combinations of workloads used to evaluate the approach. Fig. 16 presents the CDFs of error %’s for various ML models (trained using unified dataset) predicting power for different MIG combinations. To test the model’s generalizability, we evaluated workloads with a different training set from those used in Tab. II, namely BLOOM and MATMUL benchmark. As in earlier scenarios, model performance varies with workload characteristics. Notably, in MIG-specific tests, EXP2 has the lowest median comparative error amongst the three experiments. We observed that the range of metrics for EXP2 spans a very low range towards the lower end of the spectrum, approximately $\sim$95of the data points. This reflects the ability of non-identical workloads to generalize better towards fragmented ranges as opposed to generalizing well across the broad metric range spectrum.
@@ IV-B Estimation via Workload-Specific Full GPU Power Models
Finally, Fig. 18 showcases CDFs for prediction errors when models are trained using workload data specific to the MIG instances. That is, different models are used for predicting power for different MIG instances, where training data for a MIG’s model is derived from a workload that is the same as that on the MIG instance but running on a full GPU. Models constructed with MIG workload-specific metrics perform, on average, similarly to those trained on unified data. However, these workload-aware models incur high overhead due to the requirement of having individual workload-specific information available at runtime, which is not always feasible. Runtime workloads often present an opaque view, making it challenging to identify similar workloads. This highlights the advantages of a unified training approach.
@@ IV-C Partitioning Aggregate Power Using Estimates
One easy way to reduce the aggregate error to zero is to use the measured full GPU power, when available, to scale the estimated MIG powers in proportion to their estimated values (such that the sum of their scaled values equals the measured value). More specifically, letting $P_{k}^{act}$ denote the active power of the $k{\rm th}$ partition, its scaled power is given by $\frac{P_{k}^{act}}{\sum_{i}P_{i}^{act}}\times P_{GPU}^{act}$, where $P_{GPU}^{act}$ is the GPU’s measured active power. It is easy to see that the sum of all partition powers would equal the measured powered by this method. However, there is no ground truth to validate the fairness of such attribution. One test of fairness and accuracy is to see how close the active power attributed to a partition is in comparison to the power it consumes when executing alone, and whether it remains unaltered when load in other partitions changes. Note that when not scaled, the power estimated for a MIG is unaffected by the activity in other partitions, since the estimation is done independently for each partition.
@@ IV-D Estimation and Partitioning via Online MIG Power Models
Our next attempt to improve the fairness of the attribution is to construct models at run-time using the utilizations of each workload running in the MIG partitions as features, and use those models for MIG power computation. That is, instead of using the aggregate metrics for the whole GPU as features, we expand the feature set to include the metrics of each MIG separately. Thus with $n$ partitions, there would be an $n$-fold increase to the number of features. The reason for this is that, as seen already, the power characteristics of workloads can be different. Further, power consumption can depend on the data that the workloads process [28]. Hence, the differences in the workload characteristics can be made more explicit to the model by dividing each aggregate utilization metric into its MIG-level components. Although the features are MIG-level, the target variable is the whole GPU power, since that is the only available ground truth. As with previous approaches, MIG-level power is obtained in a two-step process: MIG power in the first step is obtained via the constructed model by setting the utilizations of all the MIGs except the one under consideration to 0.0. We then deduct the full GPU idle power to obtain each MIG’s active power. These active powers are then scaled using the method of Sec. IV-C.
@@
Figure 21: DCGM metrics for concurrent executions of Llama, Granite, and Bloom models on 3G, 2G, and 1G partitions, resp.
@@
Fig. 23(a) shows the plots of total GPU measured power and the total powers attributed to the three partitions as well as their sum with full GPU models, while Fig. 23(b) shows the measured GPU active power and the attributions of the same to the three partitions after scaling using the GPU power. Note that the total predicted power in the first case deviates significantly in intervals of lower utilization (e.g [200,270]), especially when the 3G partition is not loaded. This is rectified with scaling, but note that at time $\sim$65 when 3G is loaded, the attribution to 2G increases. However 3G’s power does not change when 2G’s load is reduced to 0 at time\$$\$sim$100. Similarly, at time 280, resuming 3G causes power attribution of both 1G and 2G to increase.

### v2

@@ III-B Analysis of Tensor Workloads
Having demonstrated the heterogeneity even within CUDA kernels accomplishing the same end task, we now analyze TENSOR kernels. Fig. 5 shows the distribution of key metrics of POWER, TENSO, and DRAMA across the two workloads of LLM Inference with Llama-3B model (LLAMA) and GPUBurn performing matrix multiplications, both on A100. The power consumption during GPUBurn is higher, with a broader distribution, reflecting the benchmark workload’s impact on power. Similarly, DRAMA is also higher for GPUBurn on average as opposed to LLAMA, however comprising mild peaks at $\sim$0.4. The TENSO activity on LLAMA is also lower on average as compared to GPUBurn. These differences between the two sets of distributions highlight the lack of generalization across workloads with each workload having its own set of characteristics. This imbalance renders it more difficult to accurately model the system’s behavior across all possible workload conditions. Consequently, the variability renders having a universal model towards power prediction a significant predicament, particularly with the black-box visibility into workloads in execution.
@@ III-C Hardware Heterogeneity
Dynamic variability of runtime monitored metrics on heterogeneous GPU hardware stems from disparities in monitoring granularity and resource allocation visibility across instances. Some workloads show distinct metric patterns due to high resource utilization, while others are underrepresented, leading to uneven data distributions. This imbalance complicates predictive modeling by failing to capture nuanced workload-GPU interactions. Ensuring consistent power prediction thus necessitates tailored approaches to handle workload heterogeneity dynamically at runtime.
@@ III-D Implications of Metrics
While DCGM metrics are lightweight and easily available, NVIDIA Nsight Compute (NCU) metrics provide more advanced statistical measures into the working of the GPU components. However, obtaining such metrics involve large overheads that makes it impossible to aquire them in production systems. In order to quantify the benefits provided by NCU metrics on the power consumption specifically, we train four models, Linear Regression (LR), Gradient Boosting (GB), Random Forest and XGBoost on the MATMUL workload. We chose the NCU metrics with non-trivial correlation with power for inclusion in the models. DRAM and cache access related metrics showed the highest correlation. Fig. 10 demonstrates the benefit of incorporating NCU metrics versus only relying on DCGM metrics. When using only NCU metrics, there is no significant improvement, indicating that these metrics alone may not provide substantial additional predictive power. The combined graph integrating both DCGM and NCU metrics, however shows significant improvement, suggesting that the inclusion of NCU metrics offers benefits beyond what is achieved with DCGM metrics alone, producing much lower median error. However, NCU metrics add significant overhead and are difficult to get in production systems. Furthermore, more studies are required to ascertain their value in more complex workloads using tensor cores. So we leverage lightweight DCGM metrics towards power attribution in partitioned GPUs.
@@ III-E Full GPU Power Models
TABLE II: Training Time Summary (7435 sample size)
Model Training Time (s)
Linear Regression 0.0017
Gradient Boosting 0.5671
Random Forest 1.7844
XGBoost 0.0713
Fig. 11 presents CDF (Cumulative Distribution Function) plots showing error percentages for models trained and tested across different datasets. We evaluate and compare LR, GB, RF, and XGB models. For the Granite Train/Llama Test plot, all four models demonstrate similar performance, reflecting their ability to generalize across workloads with shared characteristics. When training and testing on the same dataset (e.g., Granite Train/Granite Test or Llama Train/Llama Test), errors are low, highlighting the strong fit of the models to consistent data sources. These scenarios reveal minimal variability, demonstrating how well the models adapt to familiar patterns. In contrast, cross-validation across heterogeneous workloads, such as Granite Train/MATMUL Test, results in higher prediction errors. The largest errors occur in the Llama Train/MATMUL Test scenario, reflecting the challenges models face when predicting power consumption for workloads vastly different from the training data. LR struggles the most in these cases, with error percentages often exceeding $\sim$90%, while GB and RF fare somewhat better but still exhibit variability. Cross-validation tests further underline the importance of training data diversity. Models trained on granite and tested on GPUBurn or MATMUL show a clear increase in error percentages, particularly for LR and GB. Table II demonstrates the overhead associated with training the different types of models. The overhead runtime depicts the low computational overheads and shows that the methodology can be adapted in an online setting.
To improve generalization, mitigate errors and enhances the ability of models to handle diverse scenarios effectively, it is crucial to curate representative training sets that encompass a wide range of workload behaviors. The variability across workloads and predictor models complicates creating a unified offline model for diverse workloads. A unified model trained on combined datasets (Granite, Llama, MATMUL, GPUburn) captures both hardware behavior and workload-specific characteristics, resulting in the lowest error. This shows that aggregated data enables better prediction than isolated datasets. While LR struggles to capture non-linear GPU metric-power relationships, RF proves to be the most robust predictor, highlighting that no single methodology fits all for power prediction across heterogeneous workloads.
@@ IV Partitioning GPU Power
In this section, using GPU power characteristics learned through benchmarking, we design approaches for partitioning the aggregate GPU power among its MIG partitions. These approaches can be adapted with minimal changes to other forms of shared execution such as time-sharing and multi-process service (MPS) also. The high-level principles should be applicable in general to GPUs from other manufacturers such as Intel and AMD.
@@
Compute processors, including GPUs, are known to suffer from lack of energy proportionality[27], and consume power even in idle state when no process is active. Such power is referred to as idle power, which is generally a non-trivial amount. The idle power of NVIDIA A100-80GB is $\sim$85 W at frequencies above 1200 MHz against its peak power of $\sim$400-450 W. When a GPU executes some process, additional power is consumed, which is referred to as the GPU’s active power. This active power is dependent on the extent to which different units of the GPU are exercised. A GPU’s total power is the sum of its idle and active powers.
@@
We are now ready to present various methods for determining MIG partition power.
@@ IV-A Estimation via Generic Full GPU Power Models
This method uses the unified model (which is constructed using training data from multiple representative workloads) as described in Sec. III-E to estimate MIG instance power. First, the unified model is used to predict the power consumption of each MIG instance individually, with the instance’s normalized DCGM metrics serving as the model features. Thus, each MIG prediction is for the power that would be consumed when just that MIG’s workload executes on the GPU, and hence would include the full GPU idle power. Hence, to obtain the MIG’s active power, full GPU idle power is deducted from the total predicted power. Next, idle power for each MIG instance is computed in proportion to its size and the total size of all the active instances as described at the beginning of the section. For example, suppose that only a 2G and 3G instances are active, then they will be assigned $\frac{2}{5}^{\rm th}$ and $\frac{3}{5}^{\rm th}$ of the idle power, resp. Finally, the total power for an instance is obtained by adding its idle and active powers. The accuracy of the overall approach is evaluated by summing all instance powers and comparing the sum to the total measured GPU power.
@@
TABLE III: MIG Instance combinations in MIG paritions
Tab. III shows the various combinations of workloads used to evaluate the approach. Fig. 12 presents the CDFs of error %’s for various ML models (trained using the unified dataset) predicting power for different MIG combinations. To test the model’s generalizability, we evaluated workloads with a different training set from those used in Tab. III, namely BLOOM and MATMUL benchmark. As in earlier scenarios, model performance varies with workload characteristics. Notably, in MIG-specific tests, EXP2 has the lowest median comparative error amongst the three experiments. We observed that the range of metrics for EXP2 spans a very low range towards the lower end of the spectrum, $\sim$95of the data points. This reflects the ability of non-identical workloads to generalize better towards fragmented ranges as opposed to generalizing well across the broad metric range spectrum.
@@ IV-B Estimation via Workload-Specific Full GPU Power Models
Fig. 14 showcases CDFs for prediction errors when models are trained using workload data specific to the MIG instances. That is, different models are used for predicting power for different MIG instances, where training data for a MIG’s model is derived from a workload that is the same as that on the MIG instance but running on a full GPU. Models constructed with MIG workload-specific metrics perform, on average, similarly to those trained on unified data. However, these workload-aware models incur high overhead due to the requirement of having individual workload-specific information available at runtime, which is not always feasible. Runtime workloads often present an opaque view, making it challenging to identify similar workloads. This highlights the advantages of a unified training approach.
@@ IV-C Partitioning Aggregate Power Using Estimates
One easy way to reduce the aggregate error to zero is to use the measured full GPU power, when available, to scale the estimated MIG powers in proportion to their estimated values (such that the sum of their scaled values equals the measured value). More specifically, letting $P_{k}^{act}$ denote the active power of the $k{\rm th}$ partition, its scaled power is given by $\frac{P_{k}^{act}}{\sum_{i}P_{i}^{act}}\times P_{GPU}^{act}$, where $P_{GPU}^{act}$ is the GPU’s measured active power. It is easy to see that the sum of all partition powers would equal the measured powered by this method. One rationale/justification for the scaling (despite errors in the predicted values $P_{i}^{act}$ of the MIG partitions) is the assumption that the errors in MIG predictions are in proportion to their actual powers. However, there is no ground truth to validate the fairness of such attribution. One test of fairness and accuracy is to see how close the active power attributed to a partition is in comparison to the power it consumes when executing alone, and whether it remains unaltered when load in other partitions changes. We demonstrate that this is the case. Note that when not scaled, the power estimated for a MIG is unaffected by the activity in other partitions, since the estimation is done independently for each partition.
@@ IV-D Estimation and Partitioning via Online MIG Power Models
Our next attempt to improve the fairness of the attribution is to construct models at run-time using the utilizations of each workload running in the MIG partitions as features, and use those models for MIG power computation. As shown in Table II in Sec. III-E, model training times are negligible, especially given that MIG power/carbon reports can tolerate some delay, and hence, this approach is practically viable. With this approach, instead of using the aggregate metrics for the whole GPU as features, we expand the feature set to include the metrics of each MIG separately. Thus with $n$ partitions, there would be an $n$-fold increase to the number of features. The reason for this is that, as seen already, the power characteristics of workloads can be different. Further, power consumption can depend on the data that the workloads process [28]. Hence, the differences in the workload characteristics can be made more explicit to the model by dividing each aggregate utilization metric into its MIG-level components. Although the features are MIG-level, the target variable is the whole GPU power, since that is the only available ground truth. As with previous approaches, MIG-level power is obtained in a two-step process: MIG power in the first step is obtained via the constructed model by setting the utilizations of all the MIGs except the one under consideration to 0.0. We then deduct the full GPU idle power to obtain each MIG’s active power. These active powers are then scaled using the method of Sec. IV-C.
@@
Figure 18: DCGM metrics for concurrent executions of Llama, Granite, and Bloom models on 3G, 2G, and 1G partitions, resp. (Legend shown in the third inset applies to the first two insets, as well.)
@@
Fig. 19(a) shows the plots of total GPU measured power and the total powers attributed to the three partitions as well as their sum with full GPU models, while Fig. 19(b) shows the measured GPU active power and the attributions of the same to the three partitions after scaling using the GPU power. Note that the total predicted power in the first case deviates significantly in intervals of lower utilization (e.g [200,270]), especially when the 3G partition is not loaded. This is rectified with scaling, but note that at time $\sim$65 when 3G is loaded, the attribution to 2G increases. However 3G’s power does not change when 2G’s load is reduced to 0 at $\sim$100. Similarly, at time 280, resuming 3G causes power attribution of both 1G and 2G to increase.